. A container with glass walls has oil inside. The total length,  of the containers is 14 cm. The walls of the container have a thickness () of 25 mm and a refractive index of 1.58. The container is surrounded by air (n = 1), and the oil inside the container has a refractive index of 1.42. The angle of incidence,  is 30º. Calculate the lateral shift, x, for this container.. A container with glass walls has oil inside. The total length,  of the containers is 14 cm. The walls of the container have a thickness () of 25 mm and a refractive index of 1.58. The container is surrounded by air (n = 1), and the oil inside the container has a refractive index of 1.42. The angle of incidence,  is 30º. Calculate the lateral shift, x, for this container.  


1
Expert's answer
2022-01-13T09:20:26-0500

Solution:


Thickness of walls of container "\left(t_{1}\right)=2.5 \mathrm{~cm}"

So, Thickness of oil (which is inside the container)

"\begin{aligned}

t_{2}=l-2 t_{1} &=14 \mathrm{~cm}-2 \times 2.5 \mathrm{cm} \\

&=9 \mathrm{~cm}

\end{aligned}"

Now,

For first lateral shift of "\Delta x_{1}" (when lignt travels from air into glass wall).

"\begin{aligned}

i=30^{\circ}, & n_{1}=1, \quad n_{2}=1.58 \\

t=t_{1}=2.5 \mathrm{~cm} & \\

\text { so, } & 1 \times \sin 30^{\circ}=1.58 \times \sin r_{1} \\

& \sin r_{1}=\frac{0.5}{1.58}=0.32 \\

& r_{1}=18.45^{\circ}

\end{aligned}"

"\text { So, } \begin{aligned}

\Delta x_{1} &=2.5 \mathrm{~cm} \times \frac{\sin \left(30^{\circ}-18.45^{\circ}\right)}{\cos (18.450)} \\

&=0.528 \mathrm{~cm}

\end{aligned}"

For second latent shift of "\Delta x_{2}" (when light travels from wall into oil)

"\begin{aligned}

i_{2}=r_{1}=18.45\degree \quad n_{2}=1.58 \quad n_{3}=1.42 \\

t=t_{2}=9 \mathrm{~cm} \\

\text { so, } \quad 1.58 \times \sin (18.45)=1.42 \times \sin r_{2} \\

\Rightarrow \quad r_{2}=20.62^{\circ} \\

\text { so, } \begin{aligned}

\Delta x_{2} &=9 \mathrm{~cm} \times \frac{\sin (18.45-20.62)}{\cos (20.62)} \\

&=-0.364 \mathrm{~cm}

\end{aligned}

\end{aligned}"

"\Rightarrow negative\ \Delta x_{2}" means light is traveling from denser to rarer medium "\left(n_{2}>n_{3}\right)" . So, it bends towards normal.

For third lateral shift of "\Delta x_{3}" (when light trand from oil into wall of container)

"\begin{aligned}

&i_{3}=r_{2}=20.62^{\circ} \quad n_{4}=n_{2}=1.58 \quad n_{3}=1.42 \\

&t=t_{1}=2.5 \mathrm{~cm} \\

&\text { so, } \quad 1.42 \times \sin (20.62)=1.58 \times \sin r_{3} \\

&\Rightarrow r_{3}=18.45^{\circ}

\end{aligned}"

So,

"\begin{aligned}

\Delta x_{3} &=2.5 \mathrm{~cm} \times \frac{\sin (20.62-18.45)}{\cos 18.45} \\

&=0.1 \mathrm{~cm}

\end{aligned}"

Total lateral shift

"\begin{aligned}

\Delta x &=\Delta x_{1}+\Delta x_{2}+\Delta x_{3} \\

&=0.528 \mathrm{~cm}-0.364 \mathrm{~cm}+0.1 \mathrm{~cm}
\\ &=0.264\mathrm{~cm}

\end{aligned}"